Two cyclists are traveling in the same direction on the same bike path. One travels at 20 miles per hour and the other at 14 miles per hour. After how many hours will they be 15 miles apart?
Answer by richwmiller(17219)

20x-14x=15
6x=15
2x=5
x=2.5 hours
check
20*2.5=50
14*2.5=35
50-35=15
ok